Β 5η? - 4η Σ 6n3 +7 n=1converge or diverge
Answer #1

given 8 sn²-un = an กะ, Sก นา dt I antil ก200 | an . at 15(n+1²4(n+1) 0-26 | S(0413 จ 50 -40 6กๆ =3((วะ) -4(A+) 5 ) ค-(รก - พ=> 1 79(5-ų) n (661+ noget ans) 5) as 1-00335 -20 =) (5-0)x (6+0) = 1 So according to ratios as dt 0-2001 So Series neigther